An APC/XCB coring assembly was put together during the transit to Site 1109 in DP mode. The ship speed was ~4 nmi/hr for the 1.5 hr when the pipe was out of the water. At 1130 hr on 24 June we reduced speed to <1 nmi/hr and began lowering the drill pipe to the seafloor. A beacon was deployed at the GPS coordinates of Site 1109 (proposed site ACE-9A) at 1315 hr on 24 June 1998.


SITE 1109

Hole 1109A

The 3.5-kHz depth recorder indicated a seafloor depth of 2214 mbsl. Hole 1109A was spudded with the APC at 1630 hr on 24 June with the bit at 2210.9 mbsl. The core was virtually full, making an exact seafloor depth difficult to determine. However, there was no evidence of mud on the piston head and brown sediment inferred to be seafloor material was evident at the top; therefore, the seafloor was defined to be at 2210.9 mbsl. Recovery was 9.96 m, or 104.8%. To ensure we obtained the uppermost seafloor sediments, we decided to spud another hole.

Hole 1109B

Without offsetting the ship, we spudded Hole 1109B at 1715 hr with the bit at 2206.9 mbsl. Core 1H recovered 5.28 m; therefore, the seafloor was calculated to be at 2211.1 mbsl. This confirmed that Hole 1109A had indeed recovered nearly the entire mudline. We took a second APC core, which extended down to 14.8 mbsf. Total recovery was 15.14 m (102.3%). The cores were primarily intended to provide whole round samples for high-resolution microbiological and interstitial water (IW) chemistry studies. Hole 1109B ended when the bit was pulled clear of the seafloor at 1830 hr on 24 June.

Hole 1109C

Again without offset, we spudded Hole 1109C at 1900 hr. Core 1H was taken with the bit at 2208.9 mbsl and recovered 7.37 m, indicating a seafloor depth of 2211.0 mbsl. The APC coring continued through Core 11H (102.4 mbsf), which required 60,000 lb of overpull to get the barrel out of the formation. The APC core recovery was 107.55 m (105.0%). Cores 3H—11H were oriented with the tensor tool, and Adara temperature measurements were taken while obtaining Cores 3H, 5H, 7H, 9H, and 11H.

We began XCB coring with Core 12X, using soft formation cutting shoes through Core 37X at 346.6 mbsf. A single DVTP measurement was taken after Core 18X (169.7 mbsf). The probe tip was recovered with a very slight bend in it.

While removing the cutting shoe from Core 36X on the rig floor, the threaded connection of the shoe fractured. We suspected that the failure was caused by overheating during drilling and subsequent quenching in seawater. This prompted us to switch to hard formation cutting shoes beginning with Core 38X because the Core 37X barrel had already been deployed. We wanted to continue XCB coring before switching to RCB because of the excellent XCB core recovery (99.3% for the day). A second cutting shoe failure occurred while cutting Core 40X, and it was returned to the rig floor without the lower part of the cutting shoe. We then attempted to advance a half core to see if we could push the junk to the side of the hole, but this attempt was not successful. Erratic torque and an abraded core cutting shoe indicated that we were pushing the broken parts of the cutting shoe tip downhole, and coring operations ceased at 375.7 mbsf. Pieces of the broken cutting shoe were recovered in Core 41X once it was split. The bit was pulled clear of the seafloor at 0430 hr on 27 June, and the hole officially ended at 0800 hr when the bit cleared the rig floor. Combined APC/XCB recovery for the hole was 234.32 m (82.2%).

Hole 1109D

We then converted to an RCB coring assembly, including a 9 7/8-in bit, a mechanical bit release, and an additional stand of drill collars. After lowering the pipe to the seafloor, we spudded Hole 1109D at 1245 hr on 27 June. Seafloor was established as 2211.1 mbsl, and we drilled without coring (using an RCB center bit) to a depth of 352.8 mbsf. The drilling took 12.25 hr at an average ROP of 28.8 m/hr. We then swept the hole clean with sepiolite mud, recovered the center bit, and at 0145 hr on 28 June we started RCB coring at 352.8 mbsf. This allowed an overlap of ~23 m with the cores from Hole 1109C. We continued through Core 35R to 676.4 mbsf with few problems. Sepiolite mud pills (30 bbl) were circulated every third core as a precautionary measure, and hole conditions remained good with no fill on connections, relatively constant torque, and no overpull or drag. The formation appeared to be quite variable, which affected core recovery and penetration rates. Coring continued through Core 45R (772.9 mbsf). In this interval we began to experience some fill (2—3 m) between connections. As a result, we increased the 30-bbl mud sweeps to every core. Recovery dropped off from the 54.9% of the day before to 47.3%, and ROP dropped from 12.9 m/hr to 3.9 m/hr while in a conglomerate. No hole stability problems were encountered until after cutting Core 46R (777.6 mbsf). During pulling off bottom, the driller noted high pump pressure and high torque, indicating a portion of the hole had caved in from above. After pulling the pipe back to 762.8 mbsf with 50,000 lb overpull the hole was swept with 30 bbl of sepiolite. High pump pressure was once again noted while retrieving the core barrel and a second 30-bbl sepiolite pill was circulated. After retrieving the core barrel, a bit deplugger was deployed to clear the bit throat. Drilling parameters returned to normal, and coring proceeded through Core 51R in hard and massive diabase to a total depth of 802.0 mbsf. Recovery was exceptional in this portion of the hole (88.9%); however, penetration rates were extremely slow (1.5 to 2.0 m/hr). Although the science party wanted to penetrate below the diabase into the inferred forearc sequence below, coring was terminated because there was no indication that we were approaching the bottom of the diabase.

After sweeping the hole with 40 bbl of sepiolite mud, a short wiper trip was made in the lower (problematic) portion of the hole. The bit was back reamed to 762.8 mbsf with 40,000—50,000 lb overpull before circulation could be established. Upon running the pipe back to the bottom of the hole, the bit encountered fill at 784.8 mbsf, indicating we had lost the lowermost ~17 m of the hole. We decided not to spend the time to clean out the bottom of the hole. Another 30-bbl sepiolite mud sweep was made and two wireline runs were made to release the bit. The hole was displaced with 217 bbl of sepiolite mud and the pipe was pulled up to 98.8 mbsf in preparation for logging.

After rigging up for logging, the triple combo with the TLT was assembled and lowered down the pipe. This tool suite was able to pass several tight spots between 326.8 and 376.8 mbsf to eventually reach 782.8 mbsf. Excellent logs were recovered over nearly the entire hole up to the bottom of the pipe. However, the caliper data indicated that the tight spots between 326.8 and 376.8 mbsf had narrowed to about the tool diameter of 4 in (10.16 cm). After the triple combo was rigged down, we assembled the Formation MicroScanner (FMS) sonic tool and ran it into the hole. Because of the constrictions, this tool could not be lowered past 349.8 mbsf and we logged up from that point. Once the FMS-sonic tool was out of the hole, the logging sheaves were pulled aside and the drill string (without the bit) was lowered to 377.8 mbsf and a second run of the FMS-sonic tool was made from 775.8 mbsf up to just below the end of the drill pipe.

The final logging run was with the well seismic tool (WST) to obtain a vertical seismic profile (VSP). Because of the light weight of this tool, it had to be lowered very slowly down the pipe. We tried pumping the tool down, but this did not help and actually made the tool appear lighter because of the pump-off effect from the oil saver. Eventually, we were able to work the tool down to 460.3 mbsf (82.5 m below the end of the pipe). The VSP was conducted at nine different depths with 7—12 shots at each level. All of the logging runs obtained quality data.

After the WST was retrieved and we were rigged down from logging, we filled the hole with 80 bbl of heavy mud. The pipe was then raised to 127.8 mbsf, and the hole plugged with cement (15 bbl). The drill pipe was pulled out of the hole, clearing the seafloor at 1945 hr on 3 July. The positioning beacon was recovered at 2130 hr, and the last thruster was raised when the end of the drill string cleared the rig floor at 0015 hr on 4 July 1998. Once the rig was secured we began the transit to Site 1110 (proposed site ACE-10A). At Hole 1109D, we cored 449.2 m of section and recovered 299.87 m (67%).


To 180 Site 1110

To 180 Table of Contents